# Raster Forge: Interactive Raster Manipulation Library and GUI for Python

### v1

@@ Abstract
Current code version
v0.6.1
Permanent link to code/repository used for this code version
https://github.com/afe-oliveira/raster-forge
Legal Code License
MIT License
Code versioning system used
Git
Software code languages, tools, and services used
Python
Compilation requirements, operating environments & dependencies
Python 3.8+, PySide6, NumPy, Rasterio, Spyndex, OpenCV Python, Matplotlib
Developer documentation/manual
https://afe-oliveira.github.io/raster-forge/
Support email for questions
a22308561@alunos.ulht.pt
@@ 2.1     Library
Composites
Layers
List[NDArray, Layer]
Alpha
NDArray, Layer
Gamma
List[float], Tuple[float]
As Array
bool
[…]
Index ID
String
Index Parameters
Dictionary
Alpha
NDArray, Layer
Thresholds
List[float], Tuple[float]
Binarize
bool
As Array
bool
Slope
DEM
NDArray, Layer
Units
degrees, radians
Alpha
NDArray, Layer
As Array
bool
Aspect
DEM
NDArray, Layer
Units
degrees, radians
Alpha
NDArray, Layer
As Array
bool
@@
Alpha
NDArray, Layer
Thresholds
List[float], Tuple[float]
Invert
bool
Mask Size
3, 5
As Array
bool
Height
DTM
NDArray, Layer
DSM
NDArray, Layer
Distance
NDArray, Layer
Alpha
NDArray, Layer
As Array
bool
Fuel Map
Vegetation Coverage
NDArray, Layer
Canopy Height
NDArray, Layer
Distance
NDArray, Layer
Water Features Mask
NDArray, Layer
Artificial Structures Mask
NDArray, Layer
Fuel Models
List[int], Tuple[int]
Tree Height
float
Alpha
NDArray, Layer
As Array
bool
${}^{a}$ The multispectral indices architecture is built around the Spyndex Python library [21].
@@ 2.1.1 Containers
Array
NDArray
Bounds
Dictionary
CRS
String
Driver
String
No Data
int, float
Transform
(float, float, float, float, float, float)
Units
String
Resolution
int
Width
int
Height
int
Count
int
Mean
float
Median
float
Min
int, float
Max
int, float
Standard Deviation
float
@@
Name
Data Type
Import Layer
File Path
String Imports one band from a specified file as a layer. Applies scale to dataset if specified.
Band ID
int
Scale
int
@@
Layers
Dictionary[String: Layer]
Scale
int
@@
Name
Data Type
Import Layers
File Path
String Imports a series of bands as defined on the Config dictionary.
Config
Dictionary
Add Layer
Name
String Adds a Layer to the Layers Dictionary.
Layer
Layer, NDArray
Remove Layer
Name
String Remove a Layer from the Layers Dictionary.
Edit Layer
Old Name
String Changes the name identifier of a given Layer in the Layers dictionary.
New Name
String
@@ 3.     Impact
Raster manipulation software is used in a wide range of applications and fields, including medical imaging [23, 24], architecture [25, 26], and planetary [27, 28] and interplanetary [29, 30] remote sensing. However, as can be seen from the importance given to the geographical aspect of each raster, Raster Forge has been designed and developed with a primary focus on the field of remote sensing, and more specifically where it intersects with wildfire management.
Raster Forge was used to compute feature maps for an area of interest [31], including vegetation coverage, water, and man-made structures. The software also generated slope and aspect maps, which are crucial for calculating wildfire spread and growth. Additionally, a canopy height map was produced, providing an indication of fuel density in a specific region. Finally, Raster Forge allowed for the creation of complete fuel maps that are ready and easy to integrate into wildfire simulation environments, such as FlamMap [32], to predict the fire risk in a particular area.
Out of the particular focus of wildfire management, Raster Forge can have many other applications. The ability to extract water body maps and analyze ground moisture content can be very helpfully in hydrological modeling. Additionally, the software enables users to quickly assess the same area across various multispectral indices, which is particularly useful for gauging vegetation health, a critical aspect of agricultural endeavors.
@@ 4.     Conclusions
This work was financed by the Portuguese Agency FCT (Fundação para a Ciência e Tecnologia), in the framework of projects UIDB/00066/2020, UIDB/04111/2020 and CEECINST/00147/2018/CP1498/CT0015.

### v2

# {textblock*}

@@ Abstract
Nr. Code metadata description Please fill in this column
C1 Current code version v0.6.1
C2 Permanent link to code/repository used for this code version https://github.com/afe-oliveira/raster-forge
C3 Permanent link to Reproducible Capsule N/A
C4 Legal Code License MIT License
C5 Code versioning system used Git
C6 Software code languages, tools, and services used Python
C7 Compilation requirements, operating environments & dependencies Python 3.8+, PySide6, NumPy, Rasterio, Spyndex, OpenCV Python, Matplotlib
C8 If available Link to developer documentation/manual https://afe-oliveira.github.io/raster-forge/
C9 Support email for questions a22308561@alunos.ulht.pt
@@ 2.1     Library
Composites Layers List[NDArray, Layer]
Alpha NDArray, Layer
Gamma List[float], Tuple[float]
As Array bool
Indicesa Index ID String
Index Parameters Dictionary
Alpha NDArray, Layer
Thresholds List[float], Tuple[float]
Binarize bool
As Array bool
Slope DEM NDArray, Layer
Units degrees, radians
Alpha NDArray, Layer
As Array bool
Aspect DEM NDArray, Layer
Units degrees, radians
Alpha NDArray, Layer
As Array bool
@@
Alpha NDArray, Layer
Thresholds List[float], Tuple[float]
Invert bool
Mask Size 3, 5
As Array bool
Height DTM NDArray, Layer
DSM NDArray, Layer
Distance NDArray, Layer
Alpha NDArray, Layer
As Array bool
Fuel Map Vegetation Coverage NDArray, Layer
Canopy Height NDArray, Layer
Distance NDArray, Layer
Water Features Mask NDArray, Layer
Artificial Structures Mask NDArray, Layer
Fuel Models List[int], Tuple[int]
Tree Height float
Alpha NDArray, Layer
As Array bool
a The multispectral indices architecture is built around the Spyndex Python library [21].
@@ 2.1.1 Containers
Array NDArray
Bounds Dictionary
CRS String
Driver String
No Data int, float
Transform (float, float, float, float, float, float)
Units String
Resolution int
Width int
Height int
Count int
Mean float
Median float
Min int, float
Max int, float
Standard Deviation float
@@
Name Data Type
Import Layer File Path String Imports one band from a specified file as a layer. Applies scale to dataset if specified.
Band ID int
Scale int
@@
Layers Dictionary[String: Layer]
Scale int
@@
Name Data Type
Import Layers File Path String Imports a series of bands as defined on the Config dictionary.
Config Dictionary
Add Layer Name String Adds a Layer to the Layers Dictionary.
Layer Layer, NDArray
Remove Layer Name String Remove a Layer from the Layers Dictionary.
Edit Layer Old Name String Changes the name identifier of a given Layer in the Layers dictionary.
New Name String
@@ 3.     Impact
Raster Forge has had impact in a few projects already. The software has been used to compute terrain feature maps from Unmanned Aerial Vehicle (UAV) imagery [23]. This permitted to easily showcase results including multispectral index and vegetation coverage maps, as well as water and man-made structures masks. An early iteration of this software was also used to compute fuel type maps—along with the components necessary to build those maps, such as canopy height, vegetation coverage and soil bareness—that were integral in processing UAV data for testing/validation of a wildfire simulation model [24]. Finally, in a similar way to the previous solution, the library component of this software is also currently being used in the development of a web wildfire simulation application under the DAIS Project [25].
Expected Raster Forge short-term future impact centers around providing faster processing of large-scale aerial data for several ongoing studies. The software is being utilized to aid the construction of topographical and land-cover datasets for quickly producing terrain feature maps, such as elevation, canopy height, and vegetation density. These datasets will then be employed to train a diffusion model with the aim of generating synthetic topographical and land-cover data for wildfire model testing. In addition, Raster Forge is also being used to generate fuel maps and their sub-components (distance fields, vegetation density, man-made structures, and water feature masks, etc.) to test and validate in-development land cover and fuel type classification methods. These fuel maps are also being employed as auxiliary data to support the conceptualization and preliminary testing of a novel time and space continuous Cellular Automata (CA) wildfire model. The impact on the described studies is already being felt, as the ease-of-use and capabilities of Raster Forge are significantly expediting the data harmonization stage, thus substantially reducing the expected development times.
From the emphasis placed on the geographical aspect of each raster, Raster Forge potential impact will likely be felt in the field of remote sensing, particularly where it intersects with wildfire management. However, we believe that Raster Forge has the capability to streamline raster data processing across many other fields of study in which raster data manipulation is a fundamental concept, including medical imaging [26, 27], architecture [28, 29], and planetary [30, 31] and interplanetary [32, 33] remote sensing, broadening its potential impact.
@@ 4.     Conclusions
This work was financed by the Portuguese Agency FCT (Fundação para a Ciência e Tecnologia), in the framework of projects UIDB/00066/2020, UIDB/04111/2020, CEECINST/00002/2021/CP2788/CT0001 and CEECINST/00147/2018/CP1498/CT0015.